# Examining the influence of anisotropy on the fundamental mode of nonradial oscillation in quark stars on a complete general relativistic scheme

### v1

@@ Abstract
The anisotropic influence on the $f$-mode frequency of oscillations and dimensionless tidal deformability of strange quark matter are analyzed by employing the nonradial oscillation equations for the complete general relativity frame and tidal deformability equations, which are derived and modified from their standard form to introduce the anisotropic factor. The fluid in the compact star follows the MIT bag model with vector coupling. For the anisotropic function, we use a local anisotropy, which is regular along the whole star and is null both at the center and on the star’s surface. We show that the $f$-frequency of oscillation and dimensionless tidal deformability change considerably with the anisotropy. Finally, we investigate the correlation between the dimensionless tidal deformability of the GW170817170817 event with the anisotropy.
@@ 1 Introduction
One of the most widely accepted assumptions in the study of compact stars is that the dense fluid within these objects is isotropic. Nonetheless, theoretical evidence suggests that various phenomena may induce anisotropy in fluids under extreme conditions [24, 25, 26], see also the review articles [27, 28] for extensive discussions on anisotropic pressure in self-gravitating systems. In addition, it has been shown in recent years that the inclusion of anisotropy in relativistic star systems leads to theoretical predictions that are consistent with several observational mass-radius measurements and tidal deformability constraints [29, 30, 31, 32]. Motivated by these studies, in the present work, we investigate the impact of anisotropic pressure on the $f$-mode (fundamental mode) oscillation frequency and tidal deformability of strange stars by employing the MIT bag model with vector coupling. For this purpose, we derive the nonradial oscillation equations and the regularity conditions for the perturbation variables near the center within the framework of general relativity for anisotropic matter. These equations, together with the deformability equations, are then integrated to assess the influence of anisotropy on the fundamental oscillation modes and stellar deformability.
It is important to note that, in a similar context, the influence of anisotropy on fundamental oscillation modes was previously analyzed in [33, 34], where the authors derived the nonradial oscillation equations and presented the corresponding regularity conditions near the center. However, the equations and regularity conditions derived in our work differ from those presented in [33].
@@ 2.2 Nonradial perturbation equations
${\tilde{W}}^{\prime}=-(\ell+1)\frac{{\tilde{W}}}{r}+re^{\Lambda}\left[\frac{e^{-\Psi}{\tilde{X}}}{p_{r}+\rho+\sigma}\frac{d\rho}{dp_{r}}-\ell(\ell+1)\frac{{\tilde{V}}}{r^{2}}-\frac{{\tilde{H}}}{2}-{\tilde{K}}\right$], (2.20)
@@
$\left.\left.+\rho\right)\left[\Psi^{\prime}-\frac{1}{r}\right]-\frac{2}{r}e^{\Psi+2\Lambda}\frac{\partial\sigma}{\partial g_{11}}\right]+{\tilde{K}}\left[(p_{r}+\rho)\frac{e^{\Psi}}{2}\left[\frac{1}{r}-3\Psi^{\prime}\right]+\sigma e^{\Psi}\left[\frac{3}{r}-\Psi^{\prime}\right]\right]-{\tilde{X}}\left[\frac{\ell}{r}\right$.
$\left.+\frac{d\rho}{dp_{r}}\left[\frac{\sigma\Psi^{\prime}}{p_{r}+\rho+\sigma}+\frac{2\sigma}{r(p_{r}+\rho+\sigma)}-\frac{2}{r}\frac{\partial\sigma}{\partial\rho}\right]\right]+{\tilde{V}}\frac{e^{\Psi}}{r^{2}}\ell(\ell+1)\left[-(p_{r}+\rho)\Psi^{\prime}+\frac{2\sigma}{r}\right],$ (2.21)
@@
Here, the anisotropic profile employed depends on the fluid and spacetime variables of the form $\sigma=\sigma(p_{r},g_{11})$ with $p_{r}=p_{r}(\rho)$; see, e.g., [42, 43]. Eqs. (2.17)-(2.22) can be reduced to the fourth-order system of linear equations presented in [41] by taking ${\tilde{H}}\rightarrow-{H}$, ${\tilde{H}}_{1}\rightarrow-{H}_{1}$, ${\tilde{K}}\rightarrow-{K}$, and $\sigma=0$; review also [44]. In addition, Eqs. (2.20) and (2.21) differ from the respective equations derived in Ref. [33] in all terms where the anisotropic factor appears.
@@
giving rise to the Zerilli equation [45, 46, 47]. In Eqs. (2.23) and (2.24), $n=(\ell-1)(\ell+2)/2$ and $r^{*}$ is the so-called “tortoise" coordinate, which is related to the radial coordinate $r$ through the relation:
@@ 2.3 Tidal deformability
For strange quark stars, where the surface energy density remains nonzero, it becomes necessary to introduce a correction to $y(R)$. Because of the discontinuity present in the energy distribution, the boundary condition is adjusted as follows: [48, 49, 50, 51]
$y_{R}\longrightarrow y_{R}-\frac{4\pi R^{3}\rho_{s}}{M},$ (2.32)
with $\rho_{s}$ being the difference in energy density across the stellar surface.
@@ 3.1 Equation of state
To investigate the impact of anisotropy on the fundamental nonradial oscillation modes, we chose to use the thermodynamically consistent vector MIT bag model, where the quark interaction is mediated by the vector channel $V^{\mu}$, as we can see in Ref. [52, 53]. Thus, the Lagrangian density is given by:
$\mathcal{L}_{\rm vMIT}=\left[\bar{\psi}_{q}\big{[}\gamma^{\mu}(i\partial_{\mu}-g_{qV}V_{\mu})-m_{q}\big{]}\psi_{q}-B+\frac{1}{2}m_{V}^{2}V^{\mu}V_{\mu}\right]\Theta(\bar{\psi}_{q}\psi_{q}),$ (3.1)
where $m_{q}$ is the mass of the quark (q)=(u,d,s)(q)=(u,\,d,\,s), $B$ denotes the pressure generated by the vacuum within the “bag” that confines the quarks, $\psi_{q}$ is the Dirac quark field, and $\Theta(\bar{\psi}_{q}\psi_{q})$ is the Heaviside step function. In the case where T=0[K]T=0\,[K], we can determine the total energy density:
$\rho=\sum_{q}\epsilon_{q}+B-\frac{1}{2}m_{v}^{2}V_{0}^{2},$ (3.2)
and the total pressure,
$p_{r}=\sum_{q}\mu_{q}n_{q}-\rho.$ (3.3)
The quantity $\epsilon_{q}$ is the energy density of the quarks:
$\epsilon_{q}=\frac{N_{c}}{\pi^{2}\hbar^{3}}\int_{0}^{k_{F_{q}}}\mu_{q}k^{2}d^{3}k,$ (3.4)
where $N_{c}=3$ is the number of colors, $k_{F_{q}}$ is the Fermi momentum of the quark $q$ and $\mu_{q}$ is the chemical potential. The parameters utilized in this work are the same as presented in Ref. [52]. We use $m_{u}=m_{d}=4\,[\rm MeV]$, and $m_{s}=95\,[\rm MeV]$. We also assume a universal coupling of quarks with the vector meson, i.e., $g_{uV}=g_{dV}=g_{sV}=g_{V}$, and use some values of $G_{V}$; as defined below
$G_{V}=\bigg{(}\frac{g_{V}}{m_{V}}\bigg{)}^{2},$ (3.5)
in units of $\rm fm]^{2}$. The value of the bag constant is taken as $B=81.1\,[\rm MeV/fm^{3}]$ and $G_{V}=0.1\,[\rm fm]^{2}$.
@@ 3.2 Anisotropic profile
$\sigma=\alpha p_{r}\left(1-\frac{1}{g_{11}}\right),$ (3.6)
where $\alpha$ is a dimensionless anisotropic constant and $g_{11}=e^{2\Lambda}$. The anisotropy model (3.6) was used, e.g., to analyze the effects of the anisotropy on the radial pulsation modes of polytropic stars [54, 55] and strange quark stars [56], nonradial oscillation modes within Cowling approximation of neutron stars [42] and strange quark stars [43], magnetic field structure [57], and slowly rotating neutron stars [58, 59]. To compare our results with some results reported in the literature -see, e.g., [42, 43]- we consider $2\leq\alpha\leq 2$.
@@
$\frac{\partial\sigma}{\partial g_{11}}=\frac{\alpha p_{r}}{e^{4\Lambda}},\quad\frac{\partial\sigma}{\partial p_{r}}=\alpha\left(1-\frac{1}{g_{11}}\right),\quad{\rm and}\quad\frac{\partial\sigma}{\partial\rho}=\frac{\partial p_{r}}{\partial\rho}\frac{\partial\sigma}{\partial p_{r}}.$ (3.7)
@@ 4.1 General remarks
For tidal deformability: To examine how anisotropy influences the tidal deformability of strange stars, we numerically integrate the stellar structure equations (2.5)-(2.7) together with the tidal deformability equation (2.29) from the center of the star ($r=0$) to its surface ($r=R$). Specifically, Eqs. (2.5)-(2.7) are first solved using the fourth-order Runge–Kutta method for various choices of the anisotropy parameter $\alpha$ and central energy density $\rho_{c}$. This step provides the radial profiles for $p_{r}$, $\sigma$, $\rho$, $m$, and $\Psi$. Next, Eq. (2.7) is handled via the shooting method: an initial trial for $\Psi_{c}$ at the center is supplied, and if the resulting solution does not satisfy the boundary condition given in equation (2.7), $\Psi_{c}$ is iteratively adjusted until convergence is achieved. With the consistent $\Psi_{c}$ determined, the tidal deformability equation (2.29) is then integrated outward from $r=0$ to $r=R$, ensuring a self-consistent solution for each selected pair of $\alpha$ and $\rho_{c}$. Once the deformability function $y(r)$ is known, the value of $y(r=R)$ is determined by means of Eq. (2.3), which finally allows us to calculate the dimensionless deformability parameter (2.27).
@@ 4.2 Equilibrium configurations and frequency of oscillations
Figure 1: The total gravitational mass, in Sun’s masses, against the central energy density and as a function of the total radius are plotted on the left and right panels, respectively. In both panels, five different anisotropic parameters $\alpha$ are employed. The solid purple and black curves indicate the $95\%$ confidence intervals for the masses and radii of PSR J0030+04510030+0451 [62, 63], while the orange and gray curves show the same for PSR J0740+66200740+6620 [64, 65], both measured by the NICER collaboration. In all cases, the outer lines offer a greater likelihood of containing the true value but with lower precision, whereas the inner lines provide higher accuracy but with lower statistical coverage.
The normalized mass with Solar mass as a function of the central energy density and versus the radius is presented on the left and right panels of Fig. 1, respectively, for five different anisotropic parameters. On the right panel, the masses-radii relation reported by NICER from compact stars PSR J0030+04510030+0451 [62, 63] and PSR J0740+66200740+6620 [64, 65] are presented. The corresponding bands of the pulsars PSR J0740+66200740+6620 [66], PSR J0348+04320348+0432 [67], and PSR J1614+22301614+2230 [68], respectively marked by light yellow, light orange, and light green colors, are also shown in the mass-radius diagram. In such figure, since the central pressure becomes extremely high or low for large and small values of the anisotropic parameter, respectively, it is impossible to carry out numerical calculations with high central energy densities. Thus, we consider all equilibrium configurations where excellent numerical precision is found. Therefore, in such situations, the approach is to get as close as possible to the value corresponding to the maximum mass. In all mass against central energy density curves, the mass increases monotonically with $\rho_{c}$ until values near the maximum mass values. In all mass versus radius curves, the mass grows with the radius proportioning to $M\propto R^{3}$ until it reaches the maximum radius. At this point, all curves turn anti-clockwise, so that the masses start to increase with the diminution of the radius until it attains the maximum-mass value.
Figure 2: The $f$-mode frequency of oscillations and the normalized $\omega_{f}$ with the average density $(M/R^{3})^{0.5}$ against the total mass are presented on the left and right panels, respectively. In both panels, five different anisotropic parameters $\alpha$ are used.
The effect of anisotropic pressure on mass and radius can also be observed in Fig. 1. When examining the maximum mass and its corresponding total radius, it becomes evident that these quantities vary significantly with changes in $\alpha$. Once $\alpha$ exceeds zero, the growth of $\alpha$ requires the star to attain a higher maximum mass and a larger radius to achieve equilibrium. Moreover, the increase in mass and radius with $\sigma$ can be understood by noting the change of $\sigma$ with the parameter $\alpha$. Greater anisotropy $\sigma$ supports higher fluid pressure, enabling stars with larger masses and radii to resist collapse. Conversely, lower anisotropy leads to equilibrium states characterized by smaller mass and radius.
@@
In the left panel, all curves show that the $f$-mode frequency increases monotonically with the total mass up to the maximum-mass configuration. The influence of anisotropy on the fluid pulsation mode is also evident. From the $f$-mode frequency versus $M/M_{\odot}$ curves, in certain mass ranges, the $f$-mode frequency increases or decreases in correspondence with an increase or decrease in $\alpha$.
In the right panel, for the isotropic case ($\alpha=0$), the relationship between $\omega_{f}\sqrt{R^{3}/M}$ and $M/M_{\odot}$ is nearly linear. When $\alpha<0$, the normalized eigenfrequency $f$ decreases monotonically as $M/M_{\odot}$ increases. For $\alpha>0$, in some mass ranges, the normalized oscillation eigenfrequency $f$ decreases with increasing total mass until it reaches a minimum value of $\omega_{f}\sqrt{R^{3}/M}$. Beyond this point, the curve bends counterclockwise, and $\omega_{f}\sqrt{R^{3}/M}$ starts to increase with $M/M_{\odot}$ until the maximum-mass configuration is reached. As noted in [56], variations in the oscillation frequency are linked to changes in the radial fluid pressure induced by anisotropy.
@@
1.0-1.0 $0.5567$ $7.5531$ $0.1086$ $1.9301$ $0.9321$
0.5-0.5 $0.5950$ $7.7154$ $0.1139$ $1.9290$ $0.9279$
$400$ $0.0$ $0.6388$ $7.8925$ $0.1195$ $1.9279$ $0.9237$
+0.5+0.5 $0.6913$ $8.0870$ $0.1262$ $1.9262$ $0.9229$
+1.0+1.0 $0.7504$ $8.3013$ $0.1335$ $1.9242$ $0.9211$
1.0-1.0 $1.1918$ $9.2840$ $0.1895$ $2.1073$ $0.9437$
0.5-0.5 $1.3055$ $9.5340$ $0.2022$ $2.1032$ $0.9317$
$600$ $0.0$ $1.4386$ $9.8033$ $0.2167$ $2.0976$ $0.9252$
+0.5+0.5 $1.5964$ $10.0913$ $0.2336$ $2.0895$ $0.9135$
+1.0+1.0 $1.7802$ $10.3940$ $0.2528$ $2.0781$ $0.9020$
Table 1: The mass $M/M_{\odot}$, radius $R$, compactness $M/R$, $f$-mode oscillation frequency, and eigenfrequency normalized with the mean density $\omega_{f}\,\sqrt{R^{3}/M}$ for strange quark stars with two different central energy densities and different values of $\alpha$.
Table 1 presents the mass, radius, compactness, $f$-mode pulsation frequency, and the oscillation eigenfrequency normalized with the mean density, for compact objects for the central energy densities $400\,[\rm MeV/fm^{3}]$ and $600\,[\rm MeV/fm^{3}]$ and five values of anisotropic parameters $\alpha$. It can be noted that, when varying $\alpha$ from 1.0-1.0 to +1.0+1.0, the mass, radius, compactness, $f$-mode, and the normalized oscillation eigenfrequency change respectively in almost $35\%$, $10\%$, $\%, 3.2-3.2\%$, and $1.1\%$ for $\rho_{c}=400\,[\rm MeV/fm^{3}]$ and in $49\%$, $12\%$, $33\%$, $1.4\%$, and $4.4\%$ for $\rho_{c}=600\,[\rm MeV/fm^{3}]$. Thus, for a central energy density interval, we note that when $\alpha$ increases, the mass, radius, and compactness rise, whereas the $f$-mode frequency (and the normalized eigenfrequency $\omega_{f}$) decreases. From these results, together with those shown in Figs. 1 and 2, we see that as the star’s mass and radius increase, the oscillation mode $f$ decreases. This peculiarity has also been found in the study of compact stars in other contexts. For example, in studies of dark matter admixed with hadronic matter [69, 20] and quark matter [70, 16], in hybrid stars with hyperons and delta baryons [71], and in compact stars with crust [14].

### v2

# arXiv:2509.12438v2[gr-qc] 28 Feb 2026

@@ Abstract
The anisotropic influence on the $f$-mode frequency of oscillations and dimensionless tidal deformability of neutron stars are analyzed by employing the nonradial oscillation equations for the complete general relativity frame and tidal deformability equations, which are derived and modified from their standard form to introduce the anisotropic factor. The fluid inside the compact star obeys an equation of state constructed by matching microscopic nuclear and perturbative QCD calculations through a piecewise polytropic interpolating scheme. For the anisotropic function, we use a local anisotropy, which is regular along the whole star and is null both at the center and on the star’s surface. We show that the $f$-frequency of oscillation and dimensionless tidal deformability change considerably with the anisotropy. Finally, we investigate the correlation between the dimensionless tidal deformability of the GW170817170817 event with the anisotropy.
@@ 1 Introduction
One of the most widely accepted assumptions in the study of compact stars is that the dense fluid within these objects is isotropic. Nonetheless, theoretical evidence suggests that various phenomena may induce anisotropy in fluids under extreme conditions [74, 21, 83], see also the review articles [42, 46] for extensive discussions on anisotropic pressure in self-gravitating systems. In addition, it has been shown in recent years that the inclusion of anisotropy in relativistic star systems leads to theoretical predictions that are consistent with several observational mass-radius measurements and tidal deformability constraints [25, 26, 66, 18]. Motivated by these studies, in the present work, we investigate the impact of anisotropic pressure on the $f$-mode (fundamental mode) oscillation frequency and tidal deformability of neutron stars by using an equation of state obtained by consistently matching microscopic nuclear calculations with perturbative Quantum Chromodynamics (QCD) results, connected through a piecewise polytropic interpolation framework. For this purpose, we derive the nonradial oscillation equations and the regularity conditions for the perturbation variables near the center within the framework of general relativity for anisotropic matter. These equations, together with the deformability equations, are then integrated to assess the influence of anisotropy on the fundamental oscillation modes and stellar deformability.
It is worth noting that, in a similar context, the impact of anisotropy on quasinormal modes of neutron stars [49, 84] and on the fundamental oscillation modes of neutron stars [40, 60] and quark stars [61] has already been investigated. The nonradial oscillation equations derived in this study are consistent with those reported in [49, 40, 84], but differ from the expressions presented in [60, 61]. The difference with the last two studies occurs because, in the anisotropic case, the first law of thermodynamics needs to consider that the work done on the fluid varies with direction.
@@ 2.2 Nonradial perturbation equations
${\tilde{W}}^{\prime}=-(\ell+1)\frac{{\tilde{W}}}{r}+re^{\Lambda}\left[\frac{e^{-\Psi}{\tilde{X}}}{p_{r}+\rho}\frac{d\rho}{dp_{r}}-\ell(\ell+1)\frac{(p_{r}+\rho+\sigma)}{(p_{r}+\rho)}\frac{{\tilde{V}}}{r^{2}}-\frac{{\tilde{H}}}{2}\right$.
$\left.-\frac{2\sigma{\tilde{W}}}{e^{\Lambda}r^{2}(p_{r}+\rho)}-\frac{(p_{r}+\rho+\sigma)}{(p_{r}+\rho)}{\tilde{K}}\right$], (2.20)
@@
$\left.\left.-2\sigma\Psi^{\prime}+\frac{2\sigma^{2}}{r(p_{r}+\rho)}\right]e^{-\Lambda}\right]-{\tilde{H}}_{1}\frac{e^{\Psi}}{2}\left[(p_{r}+\rho)\left[\frac{\ell(\ell+1)}{2r}+{e^{-2\Psi}r}\omega^{2}\right]+\frac{\ell(\ell+1)\sigma}{r}\right]$
@@
$\left.\left.-2\Psi^{\prime}+\frac{2\sigma}{r(p_{r}+\rho)}\right]\right]-{\tilde{X}}\left[\frac{\ell}{r}+\frac{d\rho}{dp_{r}}\frac{2\sigma}{r(p_{r}+\rho)}-\frac{2}{r}\frac{\partial\sigma}{\partial p_{r}}\right]+{\tilde{V}}\frac{e^{\Psi}}{r^{2}}\ell(\ell+1)\left[-(p_{r}+\rho)\Psi^{\prime}\right$.
$\left.+\frac{2\sigma}{r}\right]\left[1+\frac{\sigma}{p_{r}+\rho}\right],$ (2.21)
@@
Here, the anisotropic profile employed depends on the fluid and spacetime variables of the form $\sigma=\sigma(p_{r},g_{11})$ with $p_{r}=p_{r}(\rho)$; see, e.g., [29, 15]. Eqs. (2.17)-(2.22) can be reduced to the fourth-order system of linear equations presented in [28] by taking ${\tilde{H}}\rightarrow-{H}$, ${\tilde{H}}_{1}\rightarrow-{H}_{1}$, ${\tilde{K}}\rightarrow-{K}$, and $\sigma=0$; review also [54]. In addition, Eqs. (2.20) and (2.21) differ from the respective equations derived in Ref. [60] in all terms where the anisotropic factor appears; however, these equations are in concordance with those reported in Ref. [49, 40, 84].
@@
giving rise to the Zerilli equation [85, 31, 22]. In Eqs. (2.23) and (2.24), $n=(\ell-1)(\ell+2)/2$ and $r^{*}$ is the so-called “tortoise” coordinate, which is related to the radial coordinate $r$ through the relation:
@@ 3.1 Equation of state
In the outer layers of a compact object, the composition of matter evolves from nuclei embedded in an electron sea in the outer crust to progressively more neutron-rich matter in its most inner regions. In this low density regions, the EOS has been derived from many-body calculations using phenomenological potentials, typically including both two and three nucleon interactions [12]. Recently, chiral effective field theory (CEFT) has provided a systematic framework for expanding nuclear forces at low momenta, naturally explaining the hierarchy among two-body, three-body, and weaker higher-body interactions [23, 37, 43, 41, 71, 77]. At present, microscopic calculations based on CEFT interactions allow a reliable determination of the properties of neutron star matter up to the nuclear saturation density $n_{0}$.
In the inner regions of a neutron star, at moderate high densities, the EOS of cold quark matter can be obtained in the framework of perturbative QCD. The resulting EOS depends on an unphysical parameter, namely the renormalization scale associated with the chosen scheme. Although this dependence decreases order by order in perturbation theory, it provides a practical estimate of the contributions from yet uncalculated higher-order terms and thus serves as a quantitative measure of the intrinsic theoretical uncertainty [47]. Here we use the numerical EOS derived in [48] which is well represented by a fitting function for the pressure in terms of the baryon chemical potential $\mu_{B}$ [36]. Fixing the strong coupling constant and the strange quark mass at arbitrary reference scales (using lattice and experimental data), the EOS of quark matter in $\beta$-equilibrium reads
$P_{QCD}(\mu_{B})=P_{SB}\left[c_{1}-\dfrac{a(X)}{(\mu_{B}/GeV)-b(X)}\right],$ (3.1)
with $P_{SB}=\tfrac{3}{4\pi^{2}}(\mu_{B}/3)^{4}$, $a(X)=d_{1}X^{-\nu_{1}}$ and $b(X)=d_{2}X^{-\nu_{2}}$, where $X$ is a parameter proportional to the renormalization scale of the theory (typically varied from $1$ to $4$). The parameters are $c_{1}=0.9008$, $d_{1}=0.5034$, $d_{2}=1.452$, $\nu_{1}=0.3553$ and $\nu_{2}=0.9101$. These formulae reproduce the full three-loop pressure correctly, the quark number density and the speed of sound to per cent accuracy for baryon chemical potentials up to $6$ [GeV] [47, 36].
To construct the EOS over the entire density range, we proceed as follows. For densities below $1.1n_{0}$, the CEFT EOS of [78] is used. At baryon chemical potentials above $2.6$ [GeV], where the relative uncertainty of the quark matter EOS is as large as the nuclear matter one at $n=1.1n_{0}$, the expression of Eq. (3.1) is used. Between these two regions, it is assumed that the EOS is well approximated by an interpolating polytrope built from two monotropes of the form
$P_{i}(\mu_{B})=k_{i}\left(n_{i}^{\gamma-1}+\dfrac{\gamma_{i}-1}{k_{i}\gamma_{i}}(\mu_{B}-\mu_{B,i})\right)^{\tfrac{\gamma_{i}}{\gamma_{i}-1}},$ (3.2)
where $i=1,2$, with $\mu_{B,i}$ and $n_{i}$ being the baryon chemical potential and the baryon density respectively. These functions are connected smoothly and also allowing a density jump at the matching point of the two monotropes due to a first-order phase transition. In this work we will use the representative EOS III given in [47], which satisfies the requirement that the EOS is able to support a two solar mass star (to be in agreement with recent mass determinations using pulsar timing [27, 14]).
$\begin{array}[]{c||cccc}n/n_{0}&p&\rho&\mu&c_{s}^{2}\\ \hline\cr\hline\cr 1.1&3.542&168.5&0.9775&0.15\\ 1.3&12.13&200.4&1.022&0.42\\ 1.5&34.81&234.5&1.122&0.95\\ \hline\cr 1.7&42.24&270.9&1.151&0.19\\ 1.9&49.44&308.1&1.176&0.20\\ 2.1&56.96&346.2&1.200&0.20\\ 2.3&64.79&384.9&1.222&0.20\\ 2.5&72.90&424.4&1.243&0.21\\ 2.7&81.29&464.5&1.263&0.21\\ 2.9&89.94&505.2&1.283&0.21\\ 3.1&98.84&546.5&1.301&0.22\\ 3.3&108.0&588.5&1.319&0.22\\ 3.5&117.4&631.0&1.336&0.22\\ 3.7&127.0&674.0&1.353&0.22\\ 3.9&136.8&717.5&1.369&0.23\end{array}$
Table 1: The representative equation of state III. The baryon number density $n$ is given in units of the saturation density $n_{0}=0.16\,[\rm{fm}^{3}]$, while the pressure $p$ and the energy density $\rho$ are given in $\rm{MeV/fm}^{3}]$. The solid horizontal line indicates the transition between the two monotropes [47].
@@ 3.2 Anisotropic profile
$\sigma=\alpha p_{r}\left(1-\frac{1}{g_{11}}\right)^{n},$ (3.3)
where $\alpha$ is a dimensionless anisotropic constant and $g_{11}=e^{2\Lambda}$. The anisotropy model (3.3) was used, e.g., with $n=1$, to analyze the effects of the anisotropy on the radial pulsation modes of polytropic stars [44, 17] and strange quark stars [16], nonradial oscillation modes within Cowling approximation of neutron stars [29] and strange quark stars [15], magnetic field structure [35], and slowly rotating neutron stars [73, 67] and, with $n=2$, to investigate its influence on quasinormal modes [49, 84] and on the fundamental oscillation modes of neutron stars [40]. In the present work, we consider values of $\alpha$ within the interval between 1-1 and $1$.
@@
$\frac{\partial\sigma}{\partial g_{11}}=\frac{2\alpha p_{r}}{g_{11}^{2}}\left(1-\frac{1}{g_{11}}\right),\quad\frac{\partial\sigma}{\partial p_{r}}=\alpha\left(1-\frac{1}{g_{11}}\right)^{2},\quad{\rm and}\quad\frac{\partial\sigma}{\partial\rho}=\frac{\partial p_{r}}{\partial\rho}\frac{\partial\sigma}{\partial p_{r}}.$ (3.4)
@@ 4.1 General remarks
For tidal deformability: To examine how anisotropy influences the tidal deformability of neutron stars, we numerically integrate the stellar structure equations (2.5)-(2.7) together with the tidal deformability equation (2.29) from the center of the star ($r=0$) to its surface ($r=R$). Specifically, Eqs. (2.5)-(2.7) are first solved using the fourth-order Runge–Kutta method for various choices of the anisotropy parameter $\alpha$ and central energy density $\rho_{c}$. This step provides the radial profiles for $p_{r}$, $\sigma$, $\rho$, $m$, and $\Psi$. Next, Eq. (2.7) is handled via the shooting method: an initial trial for $\Psi_{c}$ at the center is supplied, and if the resulting solution does not satisfy the boundary condition given in equation (2.7), $\Psi_{c}$ is iteratively adjusted until convergence is achieved. With the consistent $\Psi_{c}$ determined, the tidal deformability equation (2.29) is then integrated outward from $r=0$ to $r=R$, ensuring a self-consistent solution for each selected pair of $\alpha$ and $\rho_{c}$. Once the deformability function $y(r)$ is known, the value of $y(r=R)$ is determined by means of Eq. (2.3), which finally allows us to calculate the dimensionless deformability parameter (2.27).
@@ 4.2 Equilibrium configurations and frequency of oscillations
Figure 1: The total gravitational mass, in Sun’s masses, against the central energy density and as a function of the total radius are plotted on the left and right panels, respectively. In both panels, nine different anisotropic parameters $\alpha$ are employed. The solid purple and black curves indicate the $95\%$ confidence intervals for the masses and radii of PSR J0030+04510030+0451 [69, 57], while the orange and gray curves show the same for PSR J0740+66200740+6620 [70, 58], both measured by the NICER collaboration. In all cases, the outer lines offer a greater likelihood of containing the true value but with lower precision, whereas the inner lines provide higher accuracy but with lower statistical coverage.
[…]
Figure 2: The $f$-mode frequency of oscillations and the normalized $\omega_{f}$ with the average density $(M/R^{3})^{0.5}$ against the total mass are presented on the left and right panels, respectively. In both panels, nine different anisotropic parameters $\alpha$ are used.
The effect of anisotropic pressure on mass and radius can also be observed in Fig. 1. Upon examining the maximum mass and its corresponding total radius, it becomes evident that the mass varies significantly with changes in $\alpha$; however, the total radius exhibits only a small variation. Once $\alpha$ exceeds (falls) zero, the growth (diminution) of $\alpha$ requires the star to attain a higher (lower) maximum mass to achieve equilibrium. Moreover, the increase (decrease) in mass with $\sigma$ can be understood by noting the change of $\sigma$ with the parameter $\alpha$. Greater (Lesser) anisotropy $\sigma$ supports higher (lower) fluid pressure, enabling stars with larger (smaler) masses to resist collapse.
In addition, based on the curves that pass through the confidence regions reported by NICER in the mass–radius plane, it can be inferred that the model under consideration is consistent with current observational constraints. The overlap between the equilibrium configurations and the ellipses corresponding to the observed pulsars indicates that the stellar configurations predicted by the present model simultaneously reproduce the inferred masses and radii within the statistical uncertainties. This comparison allows us to constrain the permissible range of the model’s physical parameters, thereby providing additional restrictions on the degree of anisotropy and, more generally, on the equation of state of the stellar fluid.
@@
In the left panel, all curves show that the $f$-mode frequency increases monotonically with the total mass up to the maximum-mass configuration. After this point, the curve turns counterclockwise, so that the fundamental frequency of the oscillation begins to increase with the decrease in total mass. The influence of anisotropy on the fluid pulsation mode is also evident. From the $f$-mode frequency versus $M/M_{\odot}$ curves, in certain mass ranges, the $f$-mode frequency increases or decreases in correspondence with an increase or decrease in $\alpha$.
In the right panel, in all the curves presented, up to the point of maximum mass, the relationship between $\omega_{f}\sqrt{R^{3}/M}$ and $M/M_{\odot}$ produces an almost linear behavior. Beyond the maximum mass value, the curve bends clockwise and $\omega_{f}\sqrt{R^{3}/M}$ starts to increase with $M/M_{\odot}$. As noted in [16], variations in the oscillation frequency are linked to changes in the radial fluid pressure induced by anisotropy.
@@
1.0-1.0 $1.0171$ $12.792$ $0.1174$ $1.5726$ $1.2305$
0.5-0.5 $1.0277$ $12.815$ $0.1184$ $1.5902$ $1.2413$
$400$ $0.0$ $1.0386$ $12.838$ $0.1194$ $1.6082$ $1.2523$
+0.5+0.5 $1.0499$ $12.863$ $0.1205$ $1.6268$ $1.2634$
+1.0+1.0 $1.0612$ $12.888$ $0.1216$ $1.6458$ $1.2749$
1.0-1.0 $1.5498$ $11.954$ $0.1914$ $1.8863$ $1.0802$
0.5-0.5 $1.5819$ $11.995$ $0.1947$ $1.9414$ $1.1061$
$600$ $0.0$ $1.6159$ $12.038$ $0.1982$ $1.9989$ $1.1328$
+0.5+0.5 $1.6519$ $12.083$ $0.2019$ $2.0586$ $1.1602$
+1.0+1.0 $1.6902$ $12.130$ $0.2057$ $2.1187$ $1.1876$
Table 2: The mass $M/M_{\odot}$, radius $R$, compactness $M/R$, $f$-mode oscillation frequency, and eigenfrequency normalized with the mean density $\omega_{f}\,\sqrt{R^{3}/M}$ for neutron stars with two different central energy densities and different values of $\alpha$.
Table 2 presents the mass, radius, compactness, $f$-mode pulsation frequency, and the oscillation eigenfrequency normalized with the mean density, for compact objects for the central energy densities $400\,[\rm MeV/fm^{3}]$ and $600\,[\rm MeV/fm^{3}]$ and five values of anisotropic parameters $\alpha$. It can be noted that, when varying $\alpha$ from 1.0-1.0 to +1.0+1.0, the mass, radius, compactness, $f$-mode, and the normalized oscillation eigenfrequency change respectively in almost $4.41\%$, $9.60\%$, $0.42\%$, $7.32\%$, and $4.44\%$ for $\rho_{c}=400\,[\rm MeV/fm^{3}]$ and in $14\%$, $18\%$, $1.4\%$, $23\%$, and $11\%$ for $\rho_{c}=600\,[\rm MeV/fm^{3}]$. Thus, for a central energy density interval, we note that when $\alpha$ increases, the mass, radius, compactness, the $f$-mode frequency, and the normalized eigenfrequency $\omega_{f}$ rise. From these results, along with those shown in Figures 1 and 2, we see that as the star’s mass and radius increase, the oscillation mode $f$ also increases.